# Lagrangian extensions and left symmetric structures on the four-dimensional real Lie superalgebras

### v1

@@ A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    The Lie superalgebra $(2A_{1,1}+2A)^{3}_{p}:$ A Novikov structure is given by (where $\gamma\in{\mathbb{R}})$:
    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(p-\gamma_{)}e_{1}+pe_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{lll}e_{1}\cdot e_{1}=\frac{2}{p}(p^{2}\gamma+\lambda(2p^{2}-1))e_{1}+2p\lambda e_{2},&e_{1}\cdot e_{2}=-2p\lambda e_{1}+2p\gamma e_{2},\\[5.69054pt] e_{1}\cdot e_{3}=\frac{1}{p}(p^{2}\gamma+\lambda(p^{2}-1))e_{3}+\lambda e_{4},&e_{1}\cdot e_{4}=-\lambda e_{3}+p(\gamma+\lambda)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=-2p\lambda e_{1}+2p\gamma e_{2},&e_{2}\cdot e_{2}=-2p\gamma e_{1}-\frac{2}{p}(p^{2}\lambda+(2p^{2}-1)\gamma)e_{2},\\[5.69054pt] e_{2}\cdot e_{3}=-p(\lambda+\gamma)e_{3}+\gamma e_{4},&e_{2}\cdot e_{4}=-\gamma e_{3}+\frac{1}{p}((1-p^{2})\gamma-p^{2}\lambda)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\frac{2}{p}(p^{2}\gamma+(p^{2}-1)\lambda)e_{3}+2\lambda e_{4},&e_{3}\cdot e_{2}=-2p(\gamma+\lambda)e_{3}+2\gamma e_{4},\\[5.69054pt] e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=pe_{1}+pe_{2},\\[5.69054pt] e_{4}\cdot e_{1}=-2\lambda e_{3}+2p(\gamma+\lambda)e_{4}&e_{4}\cdot e_{2}=-2\gamma e_{3}-\frac{2}{p}((p^{2}-1)\gamma+p^{2}\lambda))e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=pe_{1}+pe_{2},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
@@
    The Lie superalgebra $(2A_{1,1}+2A)^{4}_{p}:$ A Novikov structure is given by (where $\gamma\in{\mathbb{R}})$:
@@
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{lll}e_{1}\cdot e_{1}=-\frac{2}{p}(p^{2}\gamma-\lambda(2p^{2}+1))e_{1}-2p\lambda e_{2},&e_{1}\cdot e_{2}=2p\lambda e_{1}-2p\gamma e_{2},\\[5.69054pt] e_{1}\cdot e_{3}=\frac{1}{p}(-p^{2}\gamma+\lambda(p^{2}+1))e_{3}+\lambda e_{4},&e_{1}\cdot e_{4}=\lambda e_{3}+p(-\gamma+\lambda)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2p\lambda e_{1}-2p\gamma e_{2},&e_{2}\cdot e_{2}=2p\gamma e_{1}-\frac{2}{p}(-p^{2}\lambda+(2p^{2}+1)\gamma)e_{2},\\[5.69054pt] e_{2}\cdot e_{3}=p(-\gamma+\lambda)e_{3}+\gamma e_{4},&e_{2}\cdot e_{4}=\gamma e_{3}+\frac{1}{p}((-1-p^{2})\gamma+p^{2}\lambda)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\frac{2}{p}(-p^{2}\gamma+(p^{2}+1)\lambda)e_{3}+2\lambda e_{4},&e_{3}\cdot e_{2}=-2p(\gamma-\lambda)e_{3}+2\gamma e_{4},\\[5.69054pt] e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=pe_{1}-pe_{2},\\[5.69054pt] e_{4}\cdot e_{1}=2\lambda e_{3}-2p(\gamma-\lambda)e_{4}&e_{4}\cdot e_{2}=2\gamma e_{3}-\frac{2}{p}((p^{2}+1)\gamma-p^{2}\lambda))e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=pe_{1}-pe_{2},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
@@
    The Lie superalgebra $(C_{1}^{1}+A):$ A Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R}):$
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(\gamma-1)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
@@ A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    The Lie superalgebra $(C_{1/2}^{1}+A):$ A Novikov structure is given by:
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2}&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4}.\end{array}$
@@
    The Lie superalgebra $(C_{-1}^{2}+A):$ A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R})$:
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{1}{2}(2+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(\gamma-2)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
@@
    The Lie superalgebra $(C^{3}+A):$ A Novikov structure is given by:
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3},&e_{1}\cdot e_{4}=e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{4}\cdot e_{1}=\gamma e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
@@
    The Lie superalgebra $(C_{0}^{5}+A):$ A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R})$:
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2}&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3}-e_{4},&e_{1}\cdot e_{4}=e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=e_{2}.\end{array}$
@@
    The Lie superalgebra $D^{1}$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in\mathbb{R})$:
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
@@
    The Lie superalgebra $D^{2}_{q}$ with $q\not=0$: A Novikov structure is given by:
[…]
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{3}\cdot e_{1}=-e_{1},&e_{3}\cdot e_{2}=-e_{1}-e_{2},&e_{3}\cdot e_{3}=\lambda e_{1}+\gamma e_{2}&e_{3}\cdot e_{4}=qe_{4}.\end{array}$
@@
    The Lie superalgebra $D^{3}_{pq}$ with $pq\not=0$: A Novikov structure is given by (where $\lambda,\mu,\gamma\in{\mathbb{R}}$:
@@
    A Balinsky-Novikov structure is given by (where $\gamma,\lambda\in\mathbb{R})$:
    $\begin{array}[]{llll}e_{3}\cdot e_{1}=-pe_{1}+e_{2},&e_{3}\cdot e_{2}=-e_{1}-pe_{2},&e_{3}\cdot e_{3}=\lambda e_{1}+\gamma e_{2}&e_{3}\cdot e_{4}=qe_{4}.\end{array}$
@@
    The Lie superalgebra $D^{11}_{pq}$ with $0<|p|\leq|q|\leq 1$: A Novikov structure is given by:
@@
    A Balinsky-Novikov structure is given by:

### v2

@@ A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    The Lie superalgebra $(2A_{1,1}+2A)^{3}_{p}:$ A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(p-\gamma)e_{1}+pe_{2},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{lll}e_{1}\cdot e_{1}=\frac{2}{p}(p^{2}\gamma+\lambda(2p^{2}-1))e_{1}+2p\lambda e_{2},&e_{1}\cdot e_{2}=-2p\lambda e_{1}+2p\gamma e_{2},\\[5.69054pt] e_{1}\cdot e_{3}=\frac{1}{p}(p^{2}\gamma+\lambda(p^{2}-1))e_{3}+\lambda e_{4},&e_{1}\cdot e_{4}=-\lambda e_{3}+p(\gamma+\lambda)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=-2p\lambda e_{1}+2p\gamma e_{2},&e_{2}\cdot e_{2}=-2p\gamma e_{1}-\frac{2}{p}(p^{2}\lambda+(2p^{2}-1)\gamma)e_{2},\\[5.69054pt] e_{2}\cdot e_{3}=-p(\lambda+\gamma)e_{3}+\gamma e_{4},&e_{2}\cdot e_{4}=-\gamma e_{3}+\frac{1}{p}((1-p^{2})\gamma-p^{2}\lambda)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\frac{2}{p}(p^{2}\gamma+(p^{2}-1)\lambda)e_{3}+2\lambda e_{4},&e_{3}\cdot e_{2}=-2p(\gamma+\lambda)e_{3}+2\gamma e_{4},\\[5.69054pt] e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=pe_{1}+pe_{2},\\[5.69054pt] e_{4}\cdot e_{1}=-2\lambda e_{3}+2p(\gamma+\lambda)e_{4},&e_{4}\cdot e_{2}=-2\gamma e_{3}-\frac{2}{p}((p^{2}-1)\gamma+p^{2}\lambda))e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=pe_{1}+pe_{2},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
@@
    The Lie superalgebra $(2A_{1,1}+2A)^{4}_{p}:$ A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{lll}e_{1}\cdot e_{1}=-\frac{2}{p}(p^{2}\gamma-\lambda(2p^{2}+1))e_{1}-2p\lambda e_{2},&e_{1}\cdot e_{2}=2p\lambda e_{1}-2p\gamma e_{2},\\[5.69054pt] e_{1}\cdot e_{3}=\frac{1}{p}(-p^{2}\gamma+\lambda(p^{2}+1))e_{3}+\lambda e_{4},&e_{1}\cdot e_{4}=\lambda e_{3}+p(-\gamma+\lambda)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=2p\lambda e_{1}-2p\gamma e_{2},&e_{2}\cdot e_{2}=2p\gamma e_{1}-\frac{2}{p}(-p^{2}\lambda+(2p^{2}+1)\gamma)e_{2},\\[5.69054pt] e_{2}\cdot e_{3}=p(-\gamma+\lambda)e_{3}+\gamma e_{4},&e_{2}\cdot e_{4}=\gamma e_{3}+\frac{1}{p}((-1-p^{2})\gamma+p^{2}\lambda)e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=\frac{2}{p}(-p^{2}\gamma+(p^{2}+1)\lambda)e_{3}+2\lambda e_{4},&e_{3}\cdot e_{2}=-2p(\gamma-\lambda)e_{3}+2\gamma e_{4},\\[5.69054pt] e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=pe_{1}-pe_{2},\\[5.69054pt] e_{4}\cdot e_{1}=2\lambda e_{3}-2p(\gamma-\lambda)e_{4},&e_{4}\cdot e_{2}=2\gamma e_{3}-\frac{2}{p}((p^{2}+1)\gamma-p^{2}\lambda))e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=pe_{1}-pe_{2},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
@@
    The Lie superalgebra $(C_{1}^{1}+A):$ A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(\gamma-1)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
@@
    The Lie superalgebra $(C_{1/2}^{1}+A):$ A Novikov structure is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4}.\end{array}$
@@
    The Lie superalgebra $(C_{-1}^{2}+A):$ A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(2+\gamma)e_{3},&e_{1}\cdot e_{4}=\frac{1}{2}(\gamma-2)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{4}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{3}=e_{2}.\end{array}$
@@
    The Lie superalgebra $(C^{3}+A):$ A Novikov structure is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3},&e_{1}\cdot e_{4}=e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{4}\cdot e_{1}=\gamma e_{4},&e_{4}\cdot e_{4}=e_{2}.\end{array}$
@@
    The Lie superalgebra $(C_{0}^{5}+A):$ A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{\gamma}{2}e_{3}-e_{4},&e_{1}\cdot e_{4}=e_{3}+\frac{\gamma}{2}e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{3}\cdot e_{3}=e_{2},&e_{4}\cdot e_{1}=\gamma e_{4},\\[5.69054pt] e_{4}\cdot e_{4}=e_{2}.\end{array}$
@@
    The Lie superalgebra $D^{1}$: A Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $D^{2}_{q}$ with $q\not=0$: A Novikov structure is given by
    $\begin{array}[]{l}e_{1}\cdot e_{3}=-qe_{1},\quad e_{2}\cdot e_{3}=-qe_{2},\quad e_{3}\cdot e_{1}=-(1+q)e_{1},\quad e_{3}\cdot e_{2}=-e_{1}-(q+1)e_{2},\\[5.69054pt] e_{3}\cdot e_{3}=-qe_{3},\quad e_{4}\cdot e_{3}=-qe_{4}.\end{array}$
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{llll}e_{3}\cdot e_{1}=-e_{1},&e_{3}\cdot e_{2}=-e_{1}-e_{2},&e_{3}\cdot e_{3}=\lambda e_{1}+\gamma e_{2},&e_{3}\cdot e_{4}=qe_{4}.\end{array}$
@@
    The Lie superalgebra $D^{3}_{pq}$ with $pq\not=0$: A Novikov structure, depending on parameters $\lambda,\mu,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{llll}e_{3}\cdot e_{1}=-pe_{1}+e_{2},&e_{3}\cdot e_{2}=-e_{1}-pe_{2},&e_{3}\cdot e_{3}=\lambda e_{1}+\gamma e_{2},&e_{3}\cdot e_{4}=qe_{4}.\end{array}$
@@
    The Lie superalgebra $D^{11}_{pq}$ with $0<|p|\leq|q|\leq 1$: A Novikov structure is given by
@@
    A Balinsky-Novikov structure is given by